OpenStudy (anonymous):

simplify [√(x+16)]-4

OpenStudy (adi3):

why dont u try first

OpenStudy (anonymous):

rewrite (srqt(x+16))^-4

OpenStudy (anonymous):

[(x+16)^1/2]^-4

OpenStudy (anonymous):

[x+16]^(1/2*-4) [x+16]^(-4/2) [x+16]^(-2)

OpenStudy (anonymous):

[sqrt(x-16)]-4 @aj01 this is the problem ☺

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I dont know how to solve it, pls help.

OpenStudy (anonymous):

i did solve it for you above

OpenStudy (anonymous):

[(x+16)^1/2]^-4 [x+16]^(1/2*-4) [x+16]^(-4/2) [x+16]^(-2)

OpenStudy (anonymous):

that -4 is not an exponent, it comes after the radical.

OpenStudy (anonymous):

i c

OpenStudy (anonymous):

[(x+16)^.5]-4 adding 16/16 to x [(16/16x+16)^.5]-4 [(16(x/16+1))^.5]-4 4(x/16+1)^.5-4 4(sqrt(x/16-1)^.5-1)

OpenStudy (anonymous):

i think the sqrt(x+16)-4 is simpler

OpenStudy (anonymous):

\[\sqrt{x+16}-4=(x+16)^{\frac{1}{2}}-4=(x+16)^{\frac{2}{4}}-2^2=((x+16)^{\frac{1}{4}})^2-2^2\] Now apply \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (jack1):

curious... is it to the power of (-4) or is it just.... minus 4...?